# Plastic tensor networks for interpretable generative modeling

### v1

@@ Abstract
A structural optimization scheme for a single-layer nonnegative adaptive tensor tree (NATT) that models a target probability distribution is proposed. The NATT scheme, by construction, has the advantage that it is interpretable as a probabilistic graphical model. We consider the NATT scheme and a recently proposed Born machine adaptive tensor tree (BMATT) optimization scheme and demonstrate their effectiveness on a variety of generative modeling tasks where the objective is to infer the hidden structure of a provided dataset. Our results show that in terms of minimizing the negative log-likelihood, the single-layer scheme has model performance comparable to the Born machine scheme, though not better. The tasks include deducing the structure of binary bitwise operations, learning the internal structure of random Bayesian networks given only visible sites, and a real-world example related to hierarchical clustering where a cladogram is constructed from mitochondrial DNA sequences. In doing so, we also show the importance of the choice of network topology and the versatility of a least-mutual information criterion in selecting a candidate structure for a tensor tree, as well as discuss aspects of these tensor tree generative models including their information content and interpretability.
1 Introduction
@@
We stress that both approaches have merits: we found that the double-layer approach, which is quantum-inspired, tends to perform better overall when it comes to identifying a correlation structure, whereas a single-layer approach that models the probability directly using nonnegative tensors allows for the classical interpretation of classical datasets. Furthermore, it is known the Born machines on tensor trees are equivalent to quantum unitary circuits, and that nonnegative tensor trees are equivalent to hidden Markov models defined on a tree [8]. This provides a guide as to where they might find a natural application as a model.
The choice of structure in TN machine learning greatly impacts the outcome of the modeling task. Depending on the nature of the dataset, some structures can be more natural candidates compared to other schemes: one-dimensional data is modeled using a chain (an MPS) [1, 2], and two-dimensional image data can be represented better with a 2D tensor tree as opposed to a chain [3]. Thus, one important aspect of generative modeling with TNs is how the geometry is to be selected: when the structure of the data is apparent, we can prescribe a good candidate. However, in general, given some data, it is hard to identify the hidden structure that best represents it. A recent method in this direction [10, 11] proposes to optimize the structure of a quantum state modeled with a tensor tree by selecting local connection geometries that minimize some information quantity defined for each bond in the network. We closely follow this principle in our proposed method and aim to demonstrate the versatility of this criterion and the general class of adaptive tensor tree (ATT) schemes on various types of data.
@@ 2 NLL optimization of a NATT
2.1 Optimization of the fused tensor
@@
$g_{\bm{P}}=T^{(a,b)}_{ijkl}-\bm{P}_{NN}(T^{(a,b)}_{ijkl}-\nabla_{T^{(a,b)}_{ijkl}}\mathcal{L})=\begin{cases}\nabla_{T^{(a,b)}_{ijkl}}\mathcal{L},\quad T^{(a,b)}_{ijkl}>\nabla_{T^{(a,b)}_{ijkl}}\mathcal{L}\\ T^{(a,b)}_{ijkl},\quad T^{(a,b)}_{ijkl}\leq\nabla_{T^{(a,b)}_{ijkl}}\mathcal{L}\end{cases}$ (7)
@@
2.2 NMF and optimization of individual tensors
@@
The computational costs of the BMATT and NATT schemes are dominated by the cost of the SVD and NMF respectively. Since the NMF procedure in our implementation is initialized using an SVD-based approach, the NATT scheme is more expensive than the BMATT scheme. Furthermore, because the NMF is iteratively computed, the cost of the NATT approach is practically determined by the convergence speed of the iterative method. We found that the number of MU iterations needed for convergence grows as size of the matrix increases. When the maximum bond dimension is $\chi_{max}=2$, the NATT scheme is almost as fast as the BMATT scheme, with the gap between them increasing as $\chi_{max}$ increases. In terms of $\chi_{max}$, the cost of the truncated SVD keeping $\chi_{max}$ singular values scales like $\mathcal{O}(\chi_{max}^{5})$ [17] (assuming a fused tensor whose four legs have bond dimension $\chi_{max}$). However, in our implementation, we computed the full SVD with cost $\mathcal{(}\chi_{max}^{6})$. MU updates minimizing the KL divergence for computing the NMF of the matricized fused tensor have cost $\mathcal{O}(\chi_{max}^{5})$ per iteration [14].
@@
$g_{\bm{P}}=T^{(a)^{\prime}}_{ijm}-\bm{P}_{NN}(T^{(a)^{\prime}}_{ijm}-\nabla_{T^{(a)^{\prime}}_{ijm}}\mathcal{L})=\begin{cases}\nabla_{T^{(a)^{\prime}}_{ijm}}\mathcal{L},\quad T^{(a)^{\prime}}_{ijm}>\nabla_{T^{(a)^{\prime}}_{ijm}}\mathcal{L}\\ T^{(a)^{\prime}}_{ijm},\quad T^{(a)^{\prime}}_{ijm}\leq\nabla_{T^{(a)^{\prime}}_{ijm}}\mathcal{L}\end{cases}$ (12)
@@
2.3 Structural optimization based on mutual information
@@
$I(A,B)\equiv\sum_{\vec{a}\in\mathcal{X}_{A}}\sum_{\vec{b}\in\mathcal{X}_{B}}P(\vec{a},\vec{b})\log\frac{P(\vec{a},\vec{b})}{P(\vec{a})P(\vec{b})}=H_{A}+H_{B}-H_{AB}$ (15)
@@
$\tilde{H}_{AB}=-\frac{1}{|X|}\sum_{\vec{x}\in X}\ln{P_{AB}(\vec{x})}$ ) (16)
$\tilde{H}_{A}=-\frac{1}{|X|}\sum_{\vec{x}\in X}\ln{P_{A}(\vec{x})}=-\frac{1}{|X|}\sum_{\vec{x}\in X}\ln{P_{AB}((\vec{x}_{A},\vec{1}_{B}))}$ ) (17)
$\tilde{H}_{B}=-\frac{1}{|X|}\sum_{\vec{x}\in X}\ln{P_{B}(\vec{x})}=-\frac{1}{|X|}\sum_{\vec{x}\in X}\ln{P_{AB}((\vec{1}_{A},\vec{x}_{B}))}$ ) (18)
@@
2.4 Information content and interpretability
One major difference between the NATT and BMATT models concerns their interpretability. Since the NATT is constrained to be nonnegative elementwise, we can directly interpret the elements as probability weights. In fact, to obtain the equivalent hidden Markov model that corresponds to the NATT, we can apply a nonnegative CP decomposition (NNCPD) on each tensor in the network, which can be done efficiently as the problem can be recast as multiple NMF subproblems for each factor matrix [20]. The NNCPD can be computed using methods generalizing the MU update [21] or using a hierarchical least squares approach [22]. Then the delta tensors correspond to nodes in the hidden Markov model, and the matrices correspond to the transition matrices in the hidden Markov model.
@@ 3 Results
$N_{samples}\equiv|X|$ $10$ $10$ $1000$ $10^{5}$ $1140$
$N_{batch}$ $10$ $10$ $1000$ $1000$ $1140$
$\eta$ $0.005$ $0.005$ $0.05$ $0.001$ $0.0005$
@@ 3.1 Random data
Figure 3: Plot of the average NLL obtained over 10 instances when training a tensor tree on 10 samples of 64 random bits. Symbols denote whether or not the structure was also optimized in the process, the colors denote the training scheme.
In Fig. 3, we plot the NLL as a function of the maximum bond dimension. For all three schemes, the NLL decreases as a function of rank, which demonstrates that the scheme works and minimizes the NLL as intended. Note that the hybrid scheme without structural optimization is omitted because it is essentially the same as the NATT without structural optimization. For the BMATT scheme, at $\chi_{max}=10$, the NLL is saturated at the bound. Without structural optimization, the NATT scheme does not consistently reach the global minimum even at $\chi_{max}=10$ since the optimization problem is generally harder due to the constraint and suboptimal structure. However, by optimizing the structure of the network, the NATT and hybrid schemes also achieve significantly lower NLL values and saturate the bound at $\chi_{max}=10$. This demonstrates that applying the least-MI principle to choose a structure greatly improves the performance of this class of tensor tree-based ML methods, and that a proper choice of structure is desirable and impacts the achievable NLL.
@@
3.2 Random data with long-range correlations
@@
Figure 5: Obtained network structures for a tensor tree trained on 10 samples of 64 random bits, where the middle 32 bits are always identical to each other and are either 0 or 1. Black lines separate the contiguous strongly-correlated portions on the left and right sides of the input string from the other inputs. For the BMATT, the bond colors denote the normalized EE, whereas for the NMF-based schemes, the bond colors correspond to the normalized MI. In the network, positions 0-15 and 48-63 are the left and right subsections respectively, and sites 16-47 are the redundant middle section.
@@
3.3 Binary bitwise operations
@@
3.4 Random binary branching Bayesian networks
@@
Figure 12: Box plots of the cluster information distance (CID) (blue) and NLL (orange) for the models trained on a Bayesian random network. We considered one instance of the problem, and for each scheme, 10 trials were done. The NLL box plots consist of those 10 samples, while the CID box plots are taken across all pairs of trials, for a total of 45 (ways to choose 2 from 10) samples. Circles indicate outliers, crosses indicate means, and the horizontal line in the boxes denote the median.
In Fig. 12, we considered a single instance of a randomly-generated branching Bayesian network, and we plot the CID and NLL across 10 trials for each scheme. The CID was computed for each pair of distinct trials, for a total of 45 data points per box plot, and the NLL box plot uses the results of each of the 10 trials. All schemes generally achieve good NLLs that are close to each other. Since the obtained CIDs are small, we conclude that the methods are all capable of successfully reproducing the target structure given only a subset of the states in the Bayesian network, and that these methods do so consistently. We again note that the NMF-based methods produce models that can be directly interpreted as a probability distribution, which is a task that may be nontrivial in the BMATT representation.
@@
To verify the correctness of the obtained model and to illustrate how one might recover transition probabilities, we plot the distribution of transition probabilities in the transition matrices for all training instances. The transition matrices were obtained by computing an approximate rank-2 NNCPD of each intermediate tensor in the model via an MU-type method [21] and normalizing each matrix row appropriately. The rank of the NNCPD for a tensor in the model corresponds to the number of hidden states at a hidden node in the associated hidden Markov model. The figure shows that the distribution of transition probabilities $p$ is strongly bimodal with modes at $p=0.8$ and $p=0.2$, which is in line with expectations given that the target Bayesian network passes its value downstream with probability $p=0.8$ and passes the complement of its value downstream with probability $p=0.2$. The variation in $p$ is due to the observation that an exact rank-2 NNCPD may not exist for a given degree-3 tensor. In the limit of a large number of data samples and an ideal optimization, we expect the tensors in the model to converge to the exact tensors, which admit an exact NNCP decomposition into the desired transition matrices.
3.5 Hierarchical real-world data: phylogenetics
@@
Finally, in Fig. 16, we show the result of Ward clustering with the same mtDNA data. The Ward clustering also produces good agreement with the currently-accepted classification, and correctly positions the red panda closer to the musteloids than to the ursoids. However, it places the small cats (Felinae) further from the big cats/panthers, separating the members of the subfamily Felidae. Like with the ATT methods, there is no assumption of an underlying evolutionary model, so the Ward clustering does not reflect changes as a function of time. As mentioned previously, with the hierarchical clustering, we do not have access to a probabilistic model that describes the relationship between the inputs, which is possible with the tensor tree methods.
4 Summary and Discussion

### v2

@@ Abstract
A structural optimization scheme for a single-layer nonnegative adaptive tensor tree (NATT) that models a target probability distribution is proposed as an alternative paradigm for generative modeling. The NATT scheme, by construction, automatically searches for a tree structure that best fits a given discrete dataset whose features serve as inputs, and has the advantage that it is interpretable as a probabilistic graphical model. We consider the NATT scheme and a recently proposed Born machine adaptive tensor tree (BMATT) optimization scheme and demonstrate their effectiveness on a variety of generative modeling tasks where the objective is to infer the hidden structure of a provided dataset. Our results show that in terms of minimizing the negative log-likelihood, the single-layer scheme has model performance comparable to the Born machine scheme, though not better. The tasks include deducing the structure of binary bitwise operations, learning the internal structure of random Bayesian networks given only visible sites, and a real-world example related to hierarchical clustering where a cladogram is constructed from mitochondrial DNA sequences. In doing so, we also show the importance of the choice of network topology and the versatility of a least-mutual information criterion in selecting a candidate structure for a tensor tree, as well as discuss aspects of these tensor tree generative models including their information content and interpretability.
I Introduction
@@
We stress that both approaches have merits: we found that the double-layer approach, which is quantum-inspired, tends to perform better overall when it comes to identifying a correlation structure, whereas a single-layer approach that models the probability directly using nonnegative tensors allows for the classical interpretation of classical datasets. Furthermore, it is known the Born machines on tensor trees are equivalent to quantum unitary circuits, and that nonnegative tensor trees are equivalent to hidden Markov models defined on a tree [8]. This provides a guide as to where they might find a natural application as a model. While there are existing approaches based on tensor sketching for generative modeling purposes [10, 11], they assume a fixed structure. There is also a sketching-based method that also adaptively determines the network topology of a graphical model given samples from the model [12], but they only apply to cases where all nodes in the model are visible nodes. The class of Markov models that our proposed method targets contain hidden variables, and all input features are associated with terminal nodes.
The choice of structure in TN machine learning greatly impacts the outcome of the modeling task. Depending on the nature of the dataset, some structures can be more natural candidates compared to other schemes: one-dimensional data is modeled using a chain (an MPS) [1, 2], and two-dimensional image data can be represented better with a 2D tensor tree as opposed to a chain [3]. Thus, one important aspect of generative modeling with TNs is how the geometry is to be selected: when the structure of the data is apparent, we can prescribe a good candidate. However, in general, given some data, it is hard to identify the hidden structure that best represents it. A recent method in this direction [13, 14] proposes to optimize the structure of a quantum state modeled with a tensor tree by selecting local connection geometries that minimize some information quantity defined for each bond in the network. We closely follow this principle in our proposed method to generate plastic structures that dynamically accommodate for input data and aim to demonstrate the versatility of this criterion and the general class of adaptive tensor tree (ATT) schemes on various types of data.
@@ II NLL optimization of a NATT
II.1 Optimization of the fused tensor
@@
$g_{\bm{P}}=T^{(a,b)}_{ijkl}-\bm{P}_{NN}(T^{(a,b)}_{ijkl}-\nabla_{T^{(a,b)}_{ijkl}}\mathcal{L})$
=$=\begin{cases}\nabla_{T^{(a,b)}_{ijkl}}\mathcal{L},\quad T^{(a,b)}_{ijkl}>\nabla_{T^{(a,b)}_{ijkl}}\mathcal{L}\\ T^{(a,b)}_{ijkl},\quad T^{(a,b)}_{ijkl}\leq\nabla_{T^{(a,b)}_{ijkl}}\mathcal{L}\end{cases}$ (7)
@@
II.2 NMF and optimization of individual tensors
@@
The computational costs of the BMATT and NATT schemes are dominated by the cost of the SVD and NMF respectively. Since the NMF procedure in our implementation is initialized using an SVD-based approach, the NATT scheme is more expensive than the BMATT scheme. Furthermore, because the NMF is iteratively computed, the cost of the NATT approach is practically determined by the convergence speed of the iterative method. We found that the number of MU iterations needed for convergence grows as size of the matrix increases. When the maximum bond dimension is $\chi_{max}=2$, the NATT scheme is almost as fast as the BMATT scheme, with the gap between them increasing as $\chi_{max}$ increases. In terms of $\chi_{max}$, the cost of the truncated SVD keeping $\chi_{max}$ singular values scales like $\mathcal{O}(\chi_{max}^{5})$ [20] (assuming a fused tensor whose four legs have bond dimension $\chi_{max}$). However, in our implementation, we computed the full SVD with cost $\mathcal{(}\chi_{max}^{6})$. MU iterations minimizing the KL divergence for computing the NMF of the matricized fused tensor have cost $\mathcal{O}(\chi_{max}^{5})$ per iteration [17].
@@
$g_{\bm{P}}=T^{(a)^{\prime}}_{ijm}-\bm{P}_{NN}(T^{(a)^{\prime}}_{ijm}-\nabla_{T^{(a)^{\prime}}_{ijm}}\mathcal{L})$
=$=\begin{cases}\nabla_{T^{(a)^{\prime}}_{ijm}}\mathcal{L},\quad T^{(a)^{\prime}}_{ijm}>\nabla_{T^{(a)^{\prime}}_{ijm}}\mathcal{L}\\ T^{(a)^{\prime}}_{ijm},\quad T^{(a)^{\prime}}_{ijm}\leq\nabla_{T^{(a)^{\prime}}_{ijm}}\mathcal{L}\end{cases}$ (12)
@@
II.3 Structural optimization based on mutual information
@@
$I(A,B)\equiv\sum_{\vec{a}\in\mathcal{X}_{A}}\sum_{\vec{b}\in\mathcal{X}_{B}}P(\vec{a},\vec{b})\log\frac{P(\vec{a},\vec{b})}{P(\vec{a})P(\vec{b})}$
$=H_{A}+H_{B}-H_{AB}$ (15)
@@
$\tilde{H}_{AB}=-\frac{1}{|X|}\sum_{\vec{x}\in X}\ln{P_{AB}(\vec{x})}$ (16)
$\tilde{H}_{A}=-\frac{1}{|X|}\sum_{\vec{x}\in X}\ln{P_{A}(\vec{x})}=-\frac{1}{|X|}\sum_{\vec{x}\in X}\ln{P_{AB}((\vec{x}_{A},\vec{1}_{B}))}$ (17)
$\tilde{H}_{B}=-\frac{1}{|X|}\sum_{\vec{x}\in X}\ln{P_{B}(\vec{x})}=-\frac{1}{|X|}\sum_{\vec{x}\in X}\ln{P_{AB}((\vec{1}_{A},\vec{x}_{B}))}$ (18)
@@
II.4 Information content and interpretability
One major difference between the NATT and BMATT models concerns their interpretability. Since the NATT is constrained to be nonnegative elementwise, we can directly interpret the elements as probability weights. In fact, to obtain the equivalent hidden Markov model that corresponds to the NATT, we can apply a nonnegative CP decomposition (NNCPD) on each tensor in the network, which can be done efficiently as the problem can be recast as multiple NMF subproblems for each factor matrix [23]. The NNCPD can be computed using methods generalizing the MU iteration [24] or using a hierarchical least squares approach [25]. Then the delta tensors correspond to nodes in the hidden Markov model, and the matrices correspond to the transition matrices in the hidden Markov model.
@@ III Results
$N_{samples}\equiv|X|$ $10$ $10$ $1000$ $10^{6}$ $1140$
$N_{batch}$ $10$ $10$ $1000$ $10^{4}$ $1140$
$\eta$ $0.005$ $0.005$ $0.05$ $0.005$ $0.0005$
@@ III.1 Random data
Figure 3: Plot of the average NLL obtained over 10 instances when training a tensor tree on 10 samples of 64 random bits. Symbols denote whether or not the structure was also optimized in the process, the colors denote the training scheme. The horizontal dotted line represents the lower bound for the NLL, which is the NLL of the dataset.
In Fig. 3, we plot the NLL as a function of the maximum bond dimension. For all three schemes, the NLL decreases as a function of rank, which demonstrates that the scheme works and minimizes the NLL as intended. Note that the hybrid scheme without structural optimization is omitted because it is essentially the same as the NATT without structural optimization. For the BMATT scheme, at $\chi_{max}=10$, the NLL approaches the bound. Without structural optimization, the NATT scheme does not consistently reach the global minimum even at $\chi_{max}=10$ since the optimization problem is generally harder due to the constraint and suboptimal structure. However, by optimizing the structure of the network, the NATT and hybrid schemes also achieve significantly lower NLL values and approach the bound at $\chi_{max}=10$. This demonstrates that applying the least-MI principle to choose a structure greatly improves the performance of this class of tensor tree-based ML methods, and that a proper choice of structure is desirable and impacts the achievable NLL.
@@
III.2 Random data with long-range correlations
@@
Figure 5: Obtained network structures for a tensor tree trained on 10 samples of 64 random bits, where the middle 32 bits are always identical to each other and are either 0 or 1. Black lines separate the contiguous strongly-correlated portions on the left and right sides of the input string from the other inputs. For the BMATT, the bond colors denote the normalized EE ($S_{EE}/\ln\chi_{max}$), whereas for the NMF-based schemes, the bond colors correspond to the normalized MI ($I/\ln\chi_{max}$). In the network, positions 0-15 and 48-63 are the left and right subsections respectively, and sites 16-47 are the redundant middle section.
@@
III.3 Binary bitwise operations
@@
III.4 Random binary branching Bayesian networks
@@
Figure 12: Box plots of the cluster information distance (CID) (blue) and NLL (orange) for the models trained on a Bayesian random network. We considered one instance of the problem, and for each scheme, 10 trials were done. The NLL box plots consist of those 10 samples, while the CID box plots are taken across all pairs of trials, for a total of 45 (ways to choose 2 from 10) samples. Circles indicate outliers, crosses indicate means, and the horizontal line in the boxes denote the median. The dashed line represents the NLL of the dataset.
In Fig. 12, we considered a single instance of a randomly-generated branching Bayesian network, and we plot the CID and NLL across 10 trials for each scheme. The CID was computed for each pair of distinct trials, for a total of 45 data points per box plot, and the NLL box plot uses the results of each of the 10 trials. All schemes generally achieve good NLLs that are close to each other. Since the obtained CIDs were all zero, all trials produced the same tree structure and we conclude that the methods are all capable of successfully reproducing the target structure given only a subset of the states in the Bayesian network, in a consistent fashion. We again note that the NMF-based methods produce models that can be directly interpreted as a probability distribution, which is a task that may be nontrivial in the BMATT representation.
@@
To verify the correctness of the obtained model and to illustrate how one might recover transition probabilities, we plot the distribution of transition probabilities in the transition matrices for all training instances. The transition matrices were obtained by computing an approximate rank-2 NNCPD of each intermediate tensor in the model via an MU-type method [24]. Computing the NNCPD for both tensors along an edge produces one matrix each along the edge. The matrices are contracted together and the result is normalized appropriately so that it can be treated as a transition matrix. The rank of the NNCPD for a tensor in the model corresponds to the number of hidden states at a hidden node in the associated hidden Markov model. The figure shows that the distribution of transition probabilities $p$ is strongly bimodal with modes at $p=0.8$ and $p=0.2$, which is in line with expectations given that the target Bayesian network passes its value downstream with probability $p=0.8$ and passes the complement of its value downstream with probability $p=0.2$. The variation in $p$ is due to the observation that an exact rank-2 NNCPD may not exist for a given degree-3 tensor. In the limit of a large number of data samples and an ideal optimization, we expect the tensors in the model to converge to the exact tensors, which admit an exact NNCP decomposition into the desired transition matrices.
III.5 Hierarchical real-world data: phylogenetics
@@
In Fig. 16, we show the result of Ward clustering with the same mtDNA data. The Ward clustering also produces good agreement with the currently-accepted classification, and correctly positions the red panda closer to the musteloids than to the ursoids. However, it places the small cats (Felinae) further from the big cats/panthers, separating the members of the subfamily Felidae. Like with the ATT methods, there is no assumption of an underlying evolutionary model, so the Ward clustering does not reflect changes as a function of time. As mentioned previously, with the hierarchical clustering, we do not have access to a probabilistic model that describes the relationship between the inputs, which is possible with the tensor tree methods.
Figure 17: Transition matrices obtained for the best tree structure for 16 species in order Carnivora, obtained using the hybrid scheme. At each terminal site, the corresponding species is identified. Along each bond, a heat map of the obtained transition matrix is provided that uses the same coloring rule as the bond colors, indicated by the bar in the right-hand side. Bond colors denote how close the transition matrix along the edge is to a permutation matrix, and this quantity is 1 if it is identical to a permutation matrix.
We can obtain an idea of the mutation probabilities between the sites in the network by computing the NNCPD of each tensor in the obtained NATT (shown in Fig. 14). Due to the ambiguity arising from relabeling of indices at each hidden site (equivalent to the insertion of permutation matrices along edges), we must consider how the obtained matrices are to be interpreted in a bit more detail.
It is generally assumed that mutation is not a common occurence, so it is expected that the transition matrices are described as having one element in each row and column that is close to one (i.e. that each matrix is close to a permutation matrix that accounts for the labeling ambiguity). Since the matrices are stochastic, the largest eigenvalue magnitude is 1, and this is reached for all eigenvalues when the transition matrix coincides with a permutation matrix. On the other hand, if the transition matrix is constant and all elements are identical, then there is one nonzero eigenvalue (which must be 1) and the remaining eigenvalues are zero. Thus, we can measure how close the matrix is to a permutation matrix by considering the smallest eigenvalue magnitude.
In Fig. 17, we observe that the bonds are all colored such that this measure is close to 1. This indicates that the matrices are all close to permutation matrices, and that the mutation probability is small. The heatmaps for each transition matrix also indicate this, and it is consistent with the expectation that the genetic distance between nearby organisms in the dataset is not significantly large.
IV Summary and Discussion